OpenStudy (anonymous):

t^2 - 3t - 10 = 0 help?

jimthompson5910 (jim_thompson5910):

Find two numbers which multiply to -10 AND add to -3

jimthompson5910 (jim_thompson5910):

Which two numbers do this?

OpenStudy (anonymous):

-5 & 2

jimthompson5910 (jim_thompson5910):

good, so the left side factors to (t-5)(t+2)

jimthompson5910 (jim_thompson5910):

therefore t^2 - 3t - 10 = 0 turns into (t-5)(t+2) = 0

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

t -5 = 0 = 5 t + 2 = 0 = -2

jimthompson5910 (jim_thompson5910):

perfect, so the solutions are t = 5 or t = -2

OpenStudy (anonymous):

okay thanks! and what about; 1/2x^2 - 3x - 8 = 0

jimthompson5910 (jim_thompson5910):

this one is a bit different because of the leading coefficient (or the first coefficient) of 1/2. But we can easily fix this by multiply EVERY term by 2 to clear out that fraction So multiply 1/2x^2 by 2 to get 2*(1/2x^2 ) = x^2 multiply -3x by 2 to get 2*(-3x) = -6x and multiply -8 by 2 to get 2*(-8) = -16 This means that 1/2x^2 - 3x - 8 = 0 turns into x^2 - 6x - 16 = 0

jimthompson5910 (jim_thompson5910):

Now we must find two numbers that both multiply to -16 AND add to -6, and they are...?

OpenStudy (anonymous):

-8 & 2

jimthompson5910 (jim_thompson5910):

good, so x^2 - 6x - 16 = 0 becomes (x-8)(x+2) = 0 which means x is??

OpenStudy (anonymous):

8 & -2?

jimthompson5910 (jim_thompson5910):

you got it, does the whole process make sense?

OpenStudy (anonymous):

yes, but then i checked with only # 8 like (8-8)(8+2) and i also got 0

jimthompson5910 (jim_thompson5910):

that's exactly right, if you check -2 you'll get the same result as well

OpenStudy (anonymous):

ohh! okk:)

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

one more thing

jimthompson5910 (jim_thompson5910):

what's that

OpenStudy (anonymous):

also my teacher gave me this formula; b^2 - 4ac

OpenStudy (anonymous):

how do i use it?

jimthompson5910 (jim_thompson5910):

that's the discriminant

jimthompson5910 (jim_thompson5910):

it will allow you to determine the number and type of solutions that a quadratic equation will have

jimthompson5910 (jim_thompson5910):

it will also determine whether you can factor or not

OpenStudy (anonymous):

can you do an example from the last problem? or; 2x^2 - 4x +1 = 0

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

ok:)

jimthompson5910 (jim_thompson5910):

let's say we want to find the number and type of solutions to x^2 - 6x - 16 = 0 In the case of x^2 - 6x - 16 , it's in the form ax^2+bx+c where a = 1, b = -6 and c = -16 Plug these values into b^2 - 4ac and simplify b^2 - 4ac (-6)^2 - 4(1)(-16) 36 - 4(1)(-16) 36 - 4(-16) 36 + 64 100 Because we got a positive number, we know that there are two real solutions (that are distinct/different) Because we got a perfect square (100 = 10^2), we know that these two solutions are rational (or fractional) solutions. Furthermore, since we got a perfect square, we know that x^2 - 6x - 16 can be factored. If we didn't get a perfect square, then we would automatically know that it couldn't be factored, which would save us time.

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

ohh okay thank you! :)

jimthompson5910 (jim_thompson5910):

sure thing :)